OpenStudy (anonymous):

Can someone please help me, I've finished most of my questions, I just need help with the substitution method. {3x + y - 2z =22 {x + 5y + z =4 {x = -3z + 6

OpenStudy (anonymous):

1) 2x - 3y + z = 5 (2) x + 3y + 8z = 22 (3) 3x - y + 2z = 12 The smart thing to notice is that if you add (1) and (2) the -3y and the +3y will eliminate the y's (1) 2x - 3y + z = 5 (2) x + 3y + 8z = 22 ----------------------- 3x + 9z = 27 and that can be divided through by 3 (4) x + 3z = 9 Now we need another equation that does not contain y. We multiply (3) by 3 and add it to (2) (3) 3x - y + 2z = 12 Multiply it by 3: 9x - 3y + 6z = 36 (2) x + 3y + 8z = 22 ----------------------- 10x +14z = 58 We can divide that through by 2 and get (5) 5x + 7z = 29 Now we put (4) and (5) together and we now have a system of only 2 equations in only 2 unknowns. (4) x + 3z = 9 (5) 5x + 7z = 29 Multiply (4) through by -5 and add to (5) -5x - 15z = -45 (5) 5x + 7z = 29 ------------------------- -8z = -16 (6) z = 2 Substitute z = 2 into (4) (4) x + 3z = 9 x + 3(2) = 9 x + 6 = 9 (7) x = 3 Substitute x=3 and z=2 into (3) (3) 3x - y + 2z = 12 3(3)- y + 2(2) = 12 9 - y + 4 = 12 13 - y = 12 -y = -1 (8) y = 1 From (7), (8), and (6), Solution = (x,y,z) = (3,1,2)

OpenStudy (anonymous):

Thank you ameen. <3

OpenStudy (anonymous):

wc